A uniform plane EM wave incident normally on a plane surface of a dielectric material is reflected with a VSWR of 3. What percentage of incident power that is reflected?


1
Expert's answer
2021-01-14T10:59:47-0500

"|\delta|^{2}=\frac{P_{reflected}}{P_{incident}}"

"P_{reflected}=|\delta|^{2}P_{incident}"

"\delta=\frac{1-|VSWR|}{1+|VSWR}=\frac{1-3}{1+3}=\frac{2}{3}=0.5"

"|\delta|^{2}=0.25"

percentage of incident power that is reflected "=0.25\times 100\%=25\%"